Assume that the risk-free rate is about 2% and the market risk premium is 8%. If you think Bank of America stock price will rise to $27 per share by the end of the year, at which time it will pay a $1 dividend, and you estimate its beta to be 1.2, what should be the fair price for BOA stock today according CAPM? the stock is currently trading at $24.16, would you buy the stock or not?
Answer #1

As per CAPM, Required Rate of Return = Risk Free Rate + (Beta*Market Risk Premium) = 2% + (1.2*8)% = 2% + 9.6% = 11.6%

Share Price After 1 year(assuming that $27 is after considering dividend) = 27+1 = $28

Therefore, Today's Fair Price as per CAPM = Price after 1 year/(1+Required Rate of Return) = 28/(1+0.116) = $25.0896

Current Price is LESS THAN Fair Price. Therefore, Stock is UNDERVALUED. Therefore, It SHOULD BE BOUGHT.